Use Euler's method with step size 0.1 to estimate y(0.2), where y(x) is the solution of the initial-value problem
y'=−5x+y^2,   y(0)=0


y(0.2)=
Answer #1

Euler Method 1) y = F(Xnry, In-1) 2) Y(X) =) 3) Xn = knyth In-It h. F (An-1, In-I) 5) n = 1, 2, 3 4) Yn Geiven :- y = -5x + yY = Yo.2) = Yy+ 0.4 :1 (-5 Hy + y Ye = Y(0-2) = 0 + 0-4 (-5X0.4+10)2] Ya = Y(0:2) = 0.1[-0.5] 92 = 9 (0.2) = -0.05 an Yn Yo =

Conclusion :- The value of y(0.2) is -0.05.

I hope it will help you .

Thanks!